The hydrogen atom at the base energy level is bombarded and stimulated with accelerated electrons with kinetic energy of 12.75 eV.
a) To which energy level is the hydrogen atom absorbing this energy?
b) Which lines and which lines are published during the return of the atom to the base energy level? Calculate the wavelengths of published lines
Answer #1

En . 13.6 ev h2 (Energy of level of Hydeogen) n-th At ground level h = 1 E,= - 13.6ev electrony = 12.75 er Energy hz .. EnergA = 2.ssev n=4 n=2 d=1240 - (486.3um 2.55 h=2 n=1 A E= 10.2 ev d = 1240 nh 121.56 10.2